# AI-Augmented Thyroid Scintigraphy for Robust Classification

### v1

@@ Abstract
Thyroid scintigraphy is a key imaging modality for diagnosing thyroid disorders. Deep learning models for thyroid scintigraphy classification often face challenges due to limited and imbalanced datasets, leading to suboptimal generalization. In this study, we investigate the effectiveness of different data augmentation techniques—Stable Diffusion (SD), Flow Matching (FM), and Conventional Augmentation (CA)—to enhance the performance of a ResNet18 classifier for thyroid condition classification. Our results showed that FM-based augmentation consistently outperforms SD-based approaches, particularly when combined with original (O) data and CA (O+FM+CA), achieving both high accuracy and fair classification across Diffuse Goiter (DG), Nodular Goiter (NG), Normal (NL), and Thyroiditis (TI) cases. The Wilcoxon statistical analysis further validated the superiority of O+FM and its variants (O+FM+CA) over SD-based augmentations in most scenarios. These findings highlight the potential of FM-based augmentation as a superior approach for generating high-quality synthetic thyroid scintigraphy images and improving model generalization in medical image classification.
@@
Thyroid Scintigraphy Image synthesis Augmentation Diffusion Stable diffusion Flow matching
@@ 1 Introduction
Thyroid diseases are among the most common endocrine disorders, affecting millions of subjects worldwide [17]. Early and accurate diagnosis is essential for effective treatment and better patient outcomes. Physicians rely on various imaging techniques, including ultrasound (US), computed tomography (CT), magnetic resonance imaging (MRI), and thyroid scintigraphy (gamma scan), along with laboratory tests, to assess thyroid conditions. Although US is widely used to evaluate nodules, its precision depends on operator’s expertise. CT and MRI are able to assess structural abnormalities but often yield nonspecific results, especially in cases like Graves’ disease [6]. By contrast, thyroid scintigraphy (using the 99mTc-pertechnetate radiopharmaceutical) provides crucial insight into the structure and function of the thyroid gland. However, interpreting these images can be subjective, time-consuming, and prone to variability among experts. These challenges highlight the need for improved, automated approaches to enhance diagnostic accuracy and efficiency [7]. Artificial intelligence (AI) has shown significant potential in the medical domain, particularly in disease diagnosis, treatment guidance, and personalized patient care [2]. However, a major challenge in developing deep learning (DL) models for medical applications is data scarcity. Large datasets are crucial for training robust models, yet collecting sufficient data can be difficult due to privacy concerns, high costs, and logistical constraints [19]. Conventional augmentation (CA) techniques, such as rotation, flipping, shifting, scaling, etc. help improve model generalization by creating variations of existing data [12]. However, these methods alone are often insufficient to fully address data limitations, highlighting the need for more advanced strategies [22]. Recent studies have investigated advanced augmentation techniques to overcome the challenge of limited medical imaging datasets. While Generative Adversarial Networks (GANs) [8] and Variational Autoencoders (VAEs) [13] have shown success, diffusion-based models [11] demonstrate superior performance in image synthesis, producing highly realistic augmented images [9, 22, 1]. This is the first study to implement a comprehensive augmentation approach on thyroid scintigraphy images using diffusion-based algorithms, including Denoising Diffusion Probabilistic Models (DDPM) [11] and Flow-Matching (FM) [15] to address the challenge of limited medical imaging data. To assess the effectiveness of the generated images, we incorporate them into the training process of a classification model and evaluate their performance on an external dataset. Contribution: Our key contributions are as follows: 1) A novel application of diffusion-based models for augmenting thyroid scintigraphy images to address data scarcity. 2) Prompt generation from physicians’ reports to maximize the use of available information in the image synthesis process. 3) A demonstration of the effectiveness of diffusion-based augmentation in enhancing classification performance for thyroid scintigraphy imaging.
2 Related-works
Zhang et al. [22] investigated SinDDM [14], a single-image denoising diffusion model, to augment lung ultrasound data. They also introduced FewDDM, an extension trained on limited samples, which outperformed single-image GANs in generating high-quality synthetic images. Augmenting with SinDDM notably improved pathology classification, especially for minority classes. Despite generating less detailed images, FewDDM surpassed SinDDM and SinGAN in downstream performance by capturing local structural variations. The study highlighted that combining synthetic and CA techniques yielded the best classification results. Hajianfar et al. [9] investigated the effectiveness of stable diffusion (SD) [18] as an advanced augmentation method in enhancing deep learning models for classifying scintigraphic thyroid images. They used reports from physicians without specific cleaning as prompts in the augmentation process. The generated images, combined with original (O) data and CA, were used to train a classifier. The results demonstrated that models trained with synthetic data achieved consistently better performance. Balla et al. [3] explored strategies to address data scarcity in musculoskeletal US for osteoarthritis detection. They used CA with diffusion-based image synthesis and the results showed that synthetic images generated through diffusion models retained anatomical fidelity and improved model generalization diagnostic accuracy, while CA sometimes hindered performance highlighting the potential of using synthetic images. Akrout et al. [1] advances data augmentation by leveraging text-to-image diffusion models to enhance a macroscopic skin disease dataset. By using text prompts, they gain fine-grained control over the image generation process. The results show that this generative augmentation approach maintains classification accuracy even when trained on a fully synthetic dataset. FM [15], as a novel, more robust, and memory-efficient method for image synthesis has shown superiority over GANs and DDPMs. While it has not been widely used in the medical domain, it has demonstrated clear advantages. In this study, we employed various strategies for advanced augmentation. Specifically, we use image masks as conditions for both DDPM and FM, and we incorporate physicians’ reports as prompts for DDPM to maximize the available information for augmentation. By leveraging FM, we aim to improve both the efficiency and quality of image synthesis, making it a key component of our approach.
@@ 3 Methodology
We aimed to find the best augmentation method to enhance the classification performance. Conventional methods [20] can generate samples that belong to completely different classes [12]. Therefore, we need a method to learn the distribution of the dataset images and then draw samples from it. GANs[8], Variational Autoencoders (VAEs) [13], DDPMs [11] and FMs [15] are some examples. Among these algorithms, DDPMs and FMs have exhibited superior performance [4]. Hence, we only consider these two approaches.
Stable Diffusion: SD is a type of Latent Diffusion Model (LDM)[18], which belongs to the DDPM family but operates in a lower-dimensional latent space to improve efficiency. LDMs first encode the image into a compact latent representation using a pre-trained VAE. The diffusion process then operates in this latent space, making it computationally efficient. For the diffusion process, there are two phases: the Forward and Reverse Processes. Let the target image be denoted as $\mathbf{x}_{0}$. In the forward process, we gradually add Gaussian noise to the sample in a Markovian manner: $\mathbf{x}_{t}=\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}+\sqrt{\beta_{t}}\mathbf{v}_{t},\quad\mathbf{v}_{t}\sim\mathcal{N}(0,I)$. The coefficients $\sqrt{1-\beta_{t}}$ and $\sqrt{\beta_{t}}$ control the transition, ensuring a gradual corruption of the data while maintaining variance stability. After a sufficient number of steps ($T$), the sample $\mathbf{x}_{T}$ follows a standard Gaussian distribution, i.e., $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$. To generate new samples, we approximate the reverse diffusion process. Starting from $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$, we iteratively sample from the conditional distribution, i.e., $\mathbf{x}_{t-1}\sim p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}).$ Since $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is intractable, we assume it follows a Gaussian distribution and train a neural network (e.g., a U-Net) to estimate the necessary parameters for sampling. By iterating this denoising process from $T$ down to 00, we obtain a generated sample $\hat{\mathbf{x}}_{0}$. Since we are using the SD model, we can incorporate additional conditioning information into the reverse process. This is achieved by modifying the reverse process to be conditional on auxiliary inputs such as text prompts or images. The new conditional distribution is given by: $\mathbf{x}_{t-1}\sim p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},C).$ where $C$ represents the selected condition, which can be $P$ (a prompt), $M$ (a mask), or $y$ (a given image). To enable conditioning, SD trains the noise prediction model $p_{\theta}$ to take both $x_{t}$ and $C$ as inputs, ensuring that the generated sample aligns with the provided condition.
Flow matching: FM provides an alternative to diffusion models by directly learning a continuous-time velocity field that defines a near-optimal transport between the source and target distributions. FM defines a straight-line (or nearly straight) transformation between samples from the data distribution and a known prior. This makes sampling more efficient compared to traditional diffusion-based approaches. Let $\mathbf{x}_{0}\sim p_{0}(x)$ and $\mathbf{x}_{1}\sim p_{1}(x)$ represent two distributions, where $\mathbf{x}_{0}$ is the source distribution (e.g., real data) and $\mathbf{x}_{1}$ is the target distribution (e.g., noise). FM constructs a continuous interpolation between these two distributions as: $\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},\quad t\in[0,1].$ This formulation defines a linear transport path from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. The goal is to learn a velocity field $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ that describes the optimal transport direction at each time step. Ideally, this velocity field should satisfy: $v_{\theta}(\mathbf{x}_{t},t)=\mathbf{x}_{1}-\mathbf{x}_{0}.$ To ensure that the learned velocity field $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ correctly follows the transport direction, we minimize the FM loss: $\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{1}}\left[|(\mathbf{x}_{1}-\mathbf{x}_{0})-v_{\theta}(\mathbf{x}_{t},t)|_{2}^{2}\right].$ During inference, novel samples can be generated by solving the learned ordinary differential equation (ODE) defined by the velocity field: $\frac{d\mathbf{x}_{t}}{dt}=v_{\theta}(\mathbf{x}_{t},t).$ This ODE governs the smooth transport from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. Unlike diffusion models, which require many discretized steps for effective denoising, FM provides a single-step or low-step approximation to recover the target distribution.
@@ 4 Experiments
Data collection: Table 1 provides an overview of the studied cases collected from nine centers using eight different imaging systems. The dataset covers a broad age range with a mean age of 44.71 ± 17.66 years. The gender distribution includes 771 males (26%) and 2,183 females (74%). The cases are classified into four categories: Diffuse Goiter (DG), Nodular Goiter (NG), Thyroiditis (TI), and Normal (NL), totaling 2,954 cases.
@@ A NA 129/303 100/203/81/48 432 ADAC GENESYS
Preprocessing: An experienced nuclear medicine physician manually segmented the thyroid region from the scintigraphy images using the manual contouring tool in ITK-Snap software [21]. Among the 2,954 images, 319 (all from center E) had a resolution of 256 × 256, while the rest were 128 × 128. Images and masks from center 5 were resampled to 128 × 128 using BSpline and nearest neighbor interpolation, respectively. Data usage for this study was approved by the research ethics board of each participating institution. In this study, we use physicians’ case reports in synthesizing images, so consistency is crucial due to varying styles and approaches across different centers. First, we analyzed all reports using GPT-4 Turbo [16] and generated questions to extract the most relevant information. These questions were then reviewed and refined by an experienced nuclear medicine physician. Finally, we used the revised questions to gather consistent information using GPT-4 Turbo and create prompts under 77 tokens to feed SD (Table 2). Additionally, for each center, an experienced technician randomly reviewed 50 cases of the generated prompts.
Table 2: Structured information extraction from thyroid scan reports
[…]
Your task is to analyze a thyroid scan report and extract structured information."
Given the following thyroid scan report: "{report}"
Answer the following questions in a structured JSON format:
"Class": "DG, NG, TI, NL",
"thyroid_function_classification": "Hypofunction / Hyperfunction / Normal / Indeterminate",
"radiotracer_uptake_pattern": "Homogeneous / Inhomogeneous / Focal / Diffuse",
"has_nodules": "Yes / No",
"multinodular_goiter": "Yes / No",
"nodule_type": "Hot / Cold / Not specified",
"thyroid_size": "Normal / Mildly enlarged / Significantly enlarged / Atrophic",
"diffuse_enlargement": "Yes / No"

### v2

@@ Abstract
Purpose: Thyroid scintigraphy plays a vital role in diagnosing a range of thyroid disorders. While deep learning classification models hold significant promise in this domain, their effectiveness is frequently compromised by limited and imbalanced datasets. This study investigates the impact of three data augmentation strategies including Stable Diffusion (SD), Flow Matching (FM), and Conventional Augmentation (CA), on enhancing the performance of a ResNet18 classifier.
Methods: Anterior thyroid scintigraphy images from 2,954 patients across nine medical centers were classified into four categories: Diffuse Goiter (DG), Nodular Goiter (NG), Normal (NL), and Thyroiditis (TI). Data augmentation was performed using various SD and FM models, resulting in 18 distinct augmentation scenarios. Each augmented dataset was used to train a ResNet18 classifier. Model performance was assessed using class-wise and average precision, recall, F1-score, AUC, and image fidelity metrics (FID and KID).
Results: FM-based augmentation outperformed all other methods, achieving the highest classification accuracy and lowest FID/KID scores, indicating both improved model generalization and realistic image synthesis. SD1, combining image and prompt inputs in the inference process, was the most effective SD variant, suggesting that physician-generated prompts provide meaningful clinical context. O+FM+CA yielded the most balanced and robust performance across all classes.
Conclusion: Integrating FM and clinically-informed SD augmentation, especially when guided by expert prompts, substantially improves thyroid scintigraphy classification. These findings highlight the importance of leveraging both structured medical input and advanced generative models for more effective training on limited datasets.
@@
Thyroid, Scintigraphy, Image synthesis, Augmentation, Diffusion, Stable diffusion, Flow matching
@@ 1 Introduction
Thyroid diseases are among the most common endocrine disorders, affecting millions of subjects worldwide [1], and their early and accurate diagnosis is essential for selecting appropriate treatments that lead to optimal patient outcomes. Physicians rely on a variety of imaging techniques, including ultrasound (US), computed tomography (CT), magnetic resonance imaging (MRI), and thyroid scintigraphy (gamma scan), along with laboratory tests, to evaluate and diagnose thyroid conditions [2].
Although US is widely used to evaluate nodular disease, its utility is dependent upon operator experience. CT and MRI can evaluate for structural characteristics, such as tracheal compression or substernal extension, but provide no functional information to assist with the diagnosis of conditions such as Graves’ disease [3]. By contrast, thyroid scintigraphy (using the 99mTc-pertechnetate radiopharmaceutical) provides crucial insight into both the structure and function of the thyroid gland. However, interpreting these images may be subjective, time-consuming, and prone to variability among experts. These challenges highlight the need for improved and objective automated approaches to enhance diagnostic accuracy and efficiency. [4].
Artificial intelligence (AI) has shown significant potential in medicine, particularly in disease diagnosis, treatment guidance, and personalized patient care [5]. However, a major challenge in developing deep learning (DL) models for medical applications is data scarcity. Large datasets are crucial for training robust models, yet collecting sufficient data can be difficult due to privacy concerns, high costs, and logistical constraints [6]. Conventional augmentation (CA) techniques, such as rotation, flipping, shifting, scaling, etc. help improve model generalization by creating variations of existing data [7]. However, these methods alone are often insufficient to fully address data limitations, highlighting the need for more advanced strategies [8].
Recent studies have investigated advanced augmentation techniques to overcome the challenge posed by limited medical imaging datasets. While Generative Adversarial Networks (GANs) [9] and Variational Autoencoders (VAEs) [10] have shown success, diffusion-based models [11] demonstrate superior performance in image synthesis, producing highly realistic augmented images [12, 8, 13].
In this study, a comprehensive augmentation approach on thyroid scintigraphy images was implemented using diffusion-based algorithms, including Denoising Diffusion Probabilistic Models (DDPM) [11] and Flow-Matching (FM) [14] to address the challenge of limited medical imaging data. To assess the effectiveness of the generated images, we incorporate them into the training process of a classification model and evaluate their performance on an external dataset.
In this work, our key contributions are as follows: 1) We explored novel application of diffusion-based models to augment thyroid scintigraphy images and address data scarcity. 2) We utilized physician reports to extract prompts used alongside input images in the Stable Diffusion models, guiding and enriching the image generation process. 3) We demonstrated that diffusion-based augmentation improves classification performance in thyroid scintigraphy imaging.
Paper architecture: The paper is organized as follows. Section 2 outlines the methods used. Section 3 covers the dataset, training setup for augmentation and classification, and evaluation metrics. Section 4 presents a detailed analysis of the results. Section 5 reviews related work and discusses our findings. Finally, Section 6 summarizes the conclusions.
@@ 2 Methodology
We aimed to find the best augmentation method to enhance the classification performance. Conventional methods [15] can generate cases that belong to completely different classes [7]. Therefore, we need a method to learn the distribution of the dataset images and then draw samples from it. GANs[9], Variational Autoencoders (VAEs) [10], DDPMs [11] and FMs [14] are some examples. Among these algorithms, DDPMs and FMs have exhibited superior performance [16]. Hence, we consider these two approaches. The flowchart of the study is presented in Figure 1.
Figure 1: Overview of the study workflow illustrating the dataset, augmentation strategies, and model training. (DG: Diffuse Goiter, NG: Nodular Goiter, NL: Normal, TI: Thyroiditis, O: Original, CA: Conventional Augmentation, SD: Stable Diffusion, FM: Flow Matching)
2.1 Stable Diffusion
SD is a type of Latent Diffusion Model (LDM) [17], which belongs to the DDPM family but operates in a lower-dimensional latent space to improve efficiency. Unlike standard DDPMs, which directly apply diffusion to raw image pixels, LDMs first encode the image into a compact latent representation using a pre-trained VAE. The diffusion process then operates in this latent space, making it computationally efficient while preserving high-quality image generation. For the diffusion process, there are two phases: the Forward Process and the Reverse Process.
Forward Process: Let the target image be denoted as $\mathbf{x}_{0}$. In the forward process, we gradually add Gaussian noise to the sample in a Markovian manner:
$\mathbf{x}_{t}=\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}+\sqrt{\beta_{t}}\mathbf{v}_{t},\quad\mathbf{v}_{t}\sim\mathcal{N}(0,I)$
The coefficients $\sqrt{1-\beta_{t}}$ and $\sqrt{\beta_{t}}$ control the transition, ensuring a gradual corruption of the data while maintaining variance stability. After a sufficient number of steps ($T$), the sample $\mathbf{x}_{T}$ follows a standard Gaussian distribution, i.e.,
$\mathbf{x}_{T}\sim\mathcal{N}(0,I)$
Reverse Process: To generate new samples, we approximate the reverse diffusion process. Starting from $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$, we iteratively sample from the conditional distribution:
$\mathbf{x}_{t-1}\sim p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$
Since $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is intractable, we assume it follows a Gaussian distribution and train a neural network (e.g., a U-Net) to estimate the necessary parameters for sampling. By iterating this denoising process from $T$ down to 0, we obtain a generated sample $\hat{\mathbf{x}}_{0}$.
Since we are using the SD model, we can incorporate additional conditioning information into the reverse process. This is achieved by modifying the reverse process to be conditional on auxiliary inputs such as text prompts or images. The new conditional distribution is given by:
$\mathbf{x}_{t-1}\sim p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},C)$
where $C$ represents the selected condition, which can be $P$ (a prompt), $M$ (a mask), or $y$ (a given image). To enable conditioning, SD trains the noise prediction model $p_{\theta}$ to take both $x_{t}$ and $C$ as inputs, ensuring that the generated sample aligns with the provided condition. This makes text-to-image and mask-to-image generation possible within the SD framework.
2.2 Flow matching
Flow Matching provides an alternative to diffusion models by directly learning a continuous-time velocity field that defines a near-optimal transport between the source and target distributions. Instead of progressively adding and then removing noise, Flow Matching defines a straight-line (or nearly straight) transformation between cases from the data distribution and a known prior. This makes sampling more efficient compared to traditional diffusion-based approaches.
Let $\mathbf{x}_{0}\sim p_{0}(x)$ and $\mathbf{x}_{1}\sim p_{1}(x)$ represent two distributions, where $\mathbf{x}_{0}$ is the source distribution (e.g., real data) and $\mathbf{x}_{1}$ is the target distribution (e.g., noise or another transformed version of the data). Flow Matching constructs a continuous interpolation between these two distributions as:
$\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},\quad t\in[0,1]$
This formulation defines a linear transport path from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. The goal is to learn a velocity field $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ that describes the optimal transport direction at each time step. Ideally, this velocity field should satisfy:
$v_{\theta}(\mathbf{x}_{t},t)=\mathbf{x}_{1}-\mathbf{x}_{0}$ (1)
To ensure that the learned velocity field $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ correctly follows the transport direction, we minimize the Flow Matching loss:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{1}}\left[|(\mathbf{x}_{1}-\mathbf{x}_{0})-v_{\theta}(\mathbf{x}_{t},t)|_{2}^{2}\right]$ (2)
This objective encourages the model to approximate the optimal transport map under a quadratic cost, ensuring that the flow remains efficient and direct.
During inference, novel cases can be generated by solving the learned ordinary differential equation (ODE) defined by the velocity field:
$\frac{d\mathbf{x}_{t}}{dt}=v_{\theta}(\mathbf{x}_{t},t)$ (3)
This ODE governs the smooth transport from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. Unlike diffusion models, which require many discretized steps for effective denoising, Flow Matching provides a single-step or low-step approximation to recover the target distribution, making it significantly more efficient in practice.
2.3 ResNet18
ResNet-18 [18] is a deep convolutional neural network (CNN) designed to enable effective feature extraction while addressing the vanishing gradient problem through residual learning. Unlike traditional CNN architectures that rely solely on stacked convolutional layers, ResNet introduces skip connections, allowing gradient flow across layers and improving convergence during training.
While attention-based architectures, such as Transformers [19], have gained popularity for complex datasets requiring long-range dependencies, our dataset does not exhibit such complexity. Instead, the patterns in our data can be effectively captured using local feature extraction mechanisms, making convolutional architectures a suitable choice. Given that ResNet-18 provides a strong balance between depth and computational efficiency, we achieve high performance without the need for more computationally expensive architectures.
ResNet-18 consists of an initial convolutional layer, followed by four residual stages, and ends with global average pooling and a fully connected layer. Each residual block applies two 3×3 convolutions with identity shortcuts, ensuring efficient feature learning. The architecture follows:
Conv(7×7,64)→MaxPool(3×3)→Residual Block×4→AvgPool→FC
@@ 3 Experiments
Table 1 provides an overview of the studied cases collected from nine centers using eight different imaging systems. The dataset covers a broad age range with a mean age of 44.71 ± 17.66 years. The gender distribution includes 771 males (26%) and 2,183 females (74%). The cases have been classified into four categories: Diffuse Goiter (DG), Nodular Goiter (NG), Thyroiditis (TI), and Normal (NL), totaling 2,954 cases. Data usage for this study was approved by the Research Ethics Committee of Kashan University of Medical Sciences (IR.KAUMS.NUHEPM.REC.1403.022), and the research was conducted in accordance with the Declaration of Helsinki.
@@ A NA 129/303 100/203/81/48 432 ADAC GENESYS
3.1 Preprocessing
An experienced nuclear medicine physician manually segmented the thyroid region from the scintigraphy images using the manual contouring tool in ITK-Snap software [20]. Among the 2,954 images, 319 (all from center E) had a resolution of 256 × 256, while the rest were 128 × 128. Images and masks from center 5 were resampled to 128 × 128 using BSpline and nearest neighbor interpolation, respectively.
In this study, we use physicians’ case reports for synthesizing images, so consistency is crucial due to varying styles and approaches across different centers. First, we analyzed all reports using GPT-4 Turbo [21] and generated questions to extract the most relevant information. These questions were then reviewed and refined by an experienced nuclear medicine physician. Finally, we used the revised questions to gather consistent information using GPT-4 Turbo and create prompts under 77 tokens to feed SD (Code Snippet 1). Additionally, for each center, an experienced technician randomly reviewed 50 cases of the generated prompts.
1prompt = f"\ You are a medical AI specialized in nuclear medicine.
2Your task is to analyze a thyroid scan report and extract structured information.
3Given the following thyroid scan report: {report}
4Answer the following questions in a structured JSON format:
5 "Class": "DG, NG, TI, NL",
6 "thyroid_function_classification": "Hypofunction / Hyperfunction / Normal / Indeterminate",
7 "radiotracer_uptake_pattern": "Homogeneous / Inhomogeneous / Focal / Diffuse",
8 "has_nodules": "Yes / No",
9 "multinodular_goiter": "Yes / No",
10 "nodule_type": "Hot / Cold / Not specified",
11 "thyroid_size": "Normal / Mildly enlarged / Significantly enlarged / Atrophic",
12 "diffuse_enlargement": "Yes / No""
Code Snippet 1: Structured information extraction from thyroid scan reports